# Automatic Speech Recognition using Advanced Deep Learning Approaches: A survey

### v1

@@ 4.1.1 Acoustic domain
Shareef et al. in [58] propose a collaborative training method for acoustic encoders in Arabic ASR systems for speech-impaired children, achieving a 10% relative accuracy improvement on phoneme alignment in the output sequence. Pioneering in recognizing impaired children’s Arabic speech. Similarly in [77], collaboratively training acoustic encoders of various sizes for on-device ASR improves efficiency and reduces redundancy. Using co-distillation with an auxiliary task, collaborative training achieves up to 11% relative WER improvement on LibriSpeech corpus.
Transducer models, in the context of ASR, map input sequences (acoustic features) to output sequences (transcriptions). Unlike traditional S2S models, transducers can handle variable-length input and output sequences more efficiently. The study [65] explores attention masking in transformer-transducer-based ASR, comparing fixed masking with chunked masking in terms of accuracy and latency. The authors claim that variable masking is the viable choice in acoustic rescoring scenarios. Similarly, to adapt the Transformer for streaming ASR, the authors in [67] employ the Transducer framework for streamable alignments. Using a unidirectional Transformer with interleaved convolution layers for audio encoding, they model future context and gradually downsample input to reduce computation cost, while limiting history context length.
Moving on, The work [66] introduces an all-in-one AM based on the Transformer architecture, combined with the CTC to ensure a sequential arrangement and utilize timing details. It addresses ASR, audio tagging (AT), and acoustic ED simultaneously. The model demonstrates superior performance, showcasing its suitability for comprehensive acoustic scene transcription.
Winata et al. [56] propose a memory-efficient Transformer architecture for end-to-end speech recognition. It significantly reduces parameters, boosting training speed by over 50% and inference time by 1.35$\times$ compared to baseline. Experiments show better generalization, lower error rates, and outperformance existing schemes without external language or acoustic models. Growing demand for on-device ASR systems prompts interest in model compression.compression.
4.1.2 Language domain
Self-attention models, such as Transformers, excel in speech recognition and reveal an important pattern. As upper self-attention layers are replaced with feed-forward layers, resembling CLDNN architecture in [48], experiments on wall street journal (WSJ) and switchboard (SWBD) datasets show no performance drop and minor gains. The novel proposed metric of attention matrix diagonality indicates increased diagonality in lower to upper encoder self-attention layers. The authors conclude that a global view appears unnecessary for training upper encoder layers in speech recognition Transformers when lower layers capture sufficient contextual information. The study conducted by Hrinchuk et al. [49] presents a proficient postprocessing model for ASR with a Transformer-based encoder-decoder architecture, initialized with the weights of pre-trained BERT model [36]. The model effectively refines ASR output, demonstrating substantial performance gains through strategies like extensive data augmentation and pretrained weight initialization. On the LibriSpeech benchmark dataset, significant reductions in WERs are observed, particularly on noisier evaluation dataset portions, outperforming baseline models and approaching the performance of Transformer-XL neural language model re-scoring with 6-gram.
Figure 8: Some end-to-end models: Basic CTC , RNN-Transducer, and attention architectures [78].
[…]
@@ 4.1.2 Language domain
CTC is an architecture and principle commonly used in S2S tasks (Figure 8), such as ASR. It enables alignment-free training by introducing a blank symbol and allowing variable-length alignments between input and output sequences. During training, the model learns to align the input sequence with the target sequence, and the blank symbol accounts for multiple possible alignments. CTC is particularly effective in tasks with variable-length outputs, making it well-suited for applications like speech recognition where the duration of spoken words may vary. Figure Figure 8 (a) illustrates the operational principles and components of CTC. This latter has been used in many ASR schemes, for example, Deng et al. [54] presents the innovative pretrained Transformer S2S ASR architecture, which integrates self-supervised pretraining techniques for comprehensive end-to-end ASR. Employing a hybrid CTC/attention model, it maximizes the potential of pretrained AM and LM. The inclusion of a CTC branch aids in the encoder’s convergence during training and considers all potential time boundaries in beam searching. The encoder is initiated with wav2vec2.0, and the introduction of a one-cross decoder (OCD) mitigates reliance on acoustic representations, enabling initialization with pretrained DistilGPT2 and overcoming the constraint of conditioning on acoustic features.
@@ 4.2.1 Acoustic domain
Schneider et al. [81] explored unsupervised pre-training for speech recognition using wav2vec model on large unlabeled audio data. The learned representations enhanced AM training with a simple CNN optimized through noise contrastive binary classification. In [82], a source filter warping data augmentation strategy is proposed to enhance the robustness of children’s speech ASR. The authors constructed an end-to-end acoustic model using the XLS-R wav2vec 2.0 model, pre-trained in a self-supervised manner on extensive cross-lingual corpora of adult speech. The work proposed in [83] introduces a multi-dialect acoustic model employing soft-parameter-sharing multi-task learning, a transductive DTL subcategory, within the Transformer architecture. Auxiliary cross-attentions aid dialect ID recognition, providing dialect information. Adaptive cross-entropy loss automatically balances multi-task learning. Experimental results demonstrate a significant reduction in error rates compared to various single- and multi-task models on multi-dialect speech recognition and dialect ID recognition tasks. Similarly, in the realm of computer vision, CNNs models like ConvNeXt have outperformed cutting-edge transformers, partly due to the integration of depthwise separable convolutions (DSC). DSC, which approximates regular convolutions, enhances the efficiency of CNNs in terms of time and memory usage without compromising accuracy—in some cases, even enhancing it. The study [84] introduces DSC into the pre-trained audio model family for audio classification on AudioSet (target task), demonstrating its advantages in balancing accuracy and model size. Xin et al. [85] introduce an audio pyramid Transformer with an attention tree structure, with four branches, to reduce computational complexity in fine-grained audio spectrogram processing. It proposes a DA transfer learning approach for weakly supervised sound event detection, a sub-field of ASR, enhancing localization performance by aligning feature distributions between frame and clip domains with a DA detection loss.
@@ 4.2.2 Language domain
The methodology is founded on the utilization of the BERT model [86], which involves pretraining language models and demonstrates improved performance across various downstream tasks. DTL approaches for language models, specifically employed in the domain of voice recognition, are referred to as LM adaptation. These approaches aim to bridge the gap between the source distribution $\mathbb{D}_{S}$ and the target distribution $\mathbb{D}_{T}$. Song et al. [87] present a novel learning-to-rescore (L2RS) approach, which relies on two main components: (i) utilizing diverse textual data from SOTA NLP models, such as BERT, and (ii) automatically determining their weights to rescore the N-best lists for ASR systems.
Recent advancements in S2S models have shown promising results for training monolingual ASR systems. The CTC and encoder-decoder models are two popular architectures for end-to-end ASR. Additionally, joint training of these architectures in a multi-task hybrid approach has been explored, demonstrating improved overall performance. For instance, the architecture illustrated in Fig. 8 (a) comprises S2S layers. The encoder network of S2S consists of a series of RNNs that generate embedding vectors, while the RNN decoder utilizes these vectors to produce final results. The RNN also benefits from prior predictions ($P_{i},i=0,\dots,n$), enhancing the accuracy of subsequent predictions. Moving on, a novel TL-based approach that enhances end-to-end speech recognition has been proposed in [88]. The novelty lies in applying DTL through multilingual training and multi-task learning at two levels. The initial stage utilizes nonnegative matrix factorization (NMF), instead of a bottleneck layer, and multilingual training for high-level feature extraction. The subsequent stage employs joint CTC-attention models on these features, where the CTC was transferred to the target attention-based model. The scheme demonstrated superior performance on TIMIT but requires testing on high-resource data. Further optimization is needed for standard end-to-end training. In addition, Integrating both AM and LM methodologies has the potential to enhance or construct an effective DTL-based ASR model, as demonstrated in [46, 89, 90, 91, 92, 93].
@@
Scheme
Based on
Speech recognition task ($\mathbb{T}_{T}$)
AM/LM
Adaptation
Result with metric
[84]
ConvNeXt-Tiny
Audio classification
AM
DA
mAP= 0.471
[85]
APT
Sound event detection
AM
DA
F1= 79.6%
[49]
BERT (Jasper)
Speech-to-text
LM
TL
WER= 14%
[54]
DistilGPT2
Improve ASR
Both
Fine-tuning
CER= 4.6%
[93]
XLRS Wave2vec
Improve ASR in low resource language
Both
Fine-tuning
5.6% WER $\downarrow$
[82]
XLRS Wave2vec
Improve ASR for children’s speech
AM
TL
WER = 4.86%
[90]
S2S
Speaker adaptation
Both
Features norm.
@@
[83]
Transformer
Multi-dialect model aids recognizing diverse speech dialects effectively
AM
Multi-task learning
Acc= 100%
[94]
S2S
Enhancing the existing multilingual S2S model.
LM
DTL
4%CER $\downarrow$ 6% WER
[95]
PaSST
Audio tagging and event detection
AM
Fine-tuning
F1= 64.85%
[81]
Wav2vec
WSJ data speech
AM
affine transform.
@@
[96]
ARoBERT
Spoken language understanding
LM
Fine-tuning
F1-score=92.56%
@@ 4.3 FL-based ASR
  • Horizontal federated fearning (HFL): In HFL, clients train a shared global model using their respective datasets, characterized by the same feature space but different sample spaces. Each client utilizes a local AI model, and their updates are aggregated by a central server without exposing raw data. The HFL training process involves: (1) initialization, (2) local training, (3) encryption of gradients, (4) secure aggregation, and (5) global model parameter updates. The objective function minimizes a global loss across all parties’ datasets [97].
  • Vertical federated learning (VFL): VFL trains models on datasets sharing the same sample space but having different feature spaces. Through entity data alignment (EDA) and encrypted model trained (EMT), VFL allows clients to cooperatively train models without sharing raw data. The training process involves the same steps as HFL [97].
@@
FL presents a paradigm shift in AI training, promoting collaboration while respecting data privacy. The working principle of HFL, and VFL is depicted on Figure 10, they cater to various data distribution scenarios, offering flexible solutions for decentralized and secure ML. The application of these FL frameworks extends across diverse domains, promising improved model accuracy and privacy preservation. The first work introducing FL in ASR is presented in [98]. The authors introduced a FL platform that is easily generalizable, incorporating hierarchical optimization and a gradient selection algorithm to enhance training time and SR performance. Guliani et al. [99] proposed a strategy to compensate non-independent and identically distributed (non-IID) data in federated training of ASR systems. The proposed strategy involved random client data sampling, which resulted in a cost-quality trade-off. Zhu et al. [100] addressed also FL-based ASR in non-IID scenarios with personalized FL. They introduced two approaches: adapting personalization layer-based FL for ASR, involving local layers for personalized models, and proposing decoupled federated learning (DecoupleFL). DecoupleFL reduces computation on clients by shifting the computation burden to the server. Additionally, it communicates secure high-level features instead of model parameters, reducing communication costs, particularly for large models. In [101], the authors proposed a client-adaptive federated training scheme to mitigate data heterogeneity when training ASR models. Nguyen et al. [102] used FL to train an ASR model based on a wav2vec 2.0 model pre-trained by self supervision. Yang et al. [103] proposed a decentralized feature extraction approach in federated learning. This approach is built upon a quantum CNN (QCNN) composed of a quantum circuit encoder for feature extraction, and an RNN based end-to-end AM. This framework takes advantage of the quantum learning progress to secure models and to avoid privacy leakage attacks. Gao et al. [104] tackled a challenging and realistic ASR federated experimental setup with clients having heterogeneous data distributions, featuring thousands of different speakers, acoustic environments, and noises. Their empirical study focused on attention-based S2S End-to-End (E2E) ASR models, evaluating three aggregation weighting strategies: standard FedAvg, loss-based aggregation, and a novel WER-based aggregation. Table 7 summarizes the most recent FL-based ASR techniques.
@@
Scheme
Based on
Speech recognition task
FL technique
Metric and result
[98]
S2S
Improve ASR
FedAvg
WER = 6%
[99]
End-to-end RNN-T
ASR on non-IDD data
FedAvg
WER=
[100]
CNN+Transformer extractor
ASR on non-IDD data
DecoupleFL
2.3- 3.4% WER $\downarrow$ compared with FedAvg
[101]
LSTM
ASR on non-IDD data
CAFT
WER = 15.13%
[102]
wav2vec 2.0
Improve ASR
FedAvg
WER= 10.92%
@@
[103]
QCNN and RNN
Improve privacy-preservation in ASR
VFL
Accuracy = 95.12%
[104]
S2S
ASR on heterogeneous data distributions
FedAvg
WER= 19.98-23.86%
[105]
S2S
ASR on private and unlabelled user data.
FedNST
WER= 22.5%
[106]
Wav2vec 2.0 and Whisper
ASR & KWS for child exploitation settings
FedAvg
WER = 11-25%
[107]
Kaldi and backoff n-gram
Improve privacy-preservation in ASR
Merging models
WER= 17.7%
[108]
TDNN
Improve privacy-preservation in ASR
Aggregation
EER = 1-2%.
[109]
Non-Streaming & Streaming Conformer
Reduce client ASR model size
Federated Dropout
6-22% $\downarrow$ Client size model; 34-3% WER $\downarrow$
4.4 RL-based ASR
@@
RL is a ML paradigm where an agent learns optimal decision-making by interacting with an environment. The agent receives feedback in the form of rewards or penalties, adapting its behavior to maximize cumulative reward over time through a trial-and-error process. reinforcement learning (RL) involves several key concepts, as defined in the following terms: Environment model, serving as a representation of contextual dynamics; State (s), denoting the current situation perceived by the agent; Observation (o), a subset of the state directly perceived by the agent; Action (a), the decision made by the agent in response to the environment; Policy ($\pi$), describing how the agent converts environmental conditions into actions; Agent, the entity making decisions based on current states and past experiences; Reward, numerical values assigned by the environment to the agent based on state-action interactions. Figure 11 illustrate the principle of RL.
Figure 11: RL principle.
@@
Markov Decision Process (MDP) is a fundamental framework for dynamic and stochastic decision-making, characterized by state space $S$, action space $\mathbb{A}$, transition probabilities $\mathbb{P}$, and a reward function $R$. The primary objective in an MDP is to identify an optimal policy $\pi^{*}$ maximizing the expected discounted total reward over time, expressed as $\pi^{*}=\max_{\pi}\mathbb{E}_{\pi}[\sum_{t=0}^{T}\gamma^{t}r_{t}(s_{t},\pi(s_{t}))]$, where $\gamma$ is the discount factor. MDPs find extensive applications in addressing uncertainties in intelligent systems within dynamic wireless environments, including spectrum management, cognitive radios, and wireless security.
@@
In the field of ASR, RL has primarily been proposed to tackle discrepancies between training and testing phases. Two main discrepancies leading to potential performance deterioration have been identified: 1) The conventional use of the cross-entropy criterion maximizes log-likelihood during training, while performance is assessed by WER, not log-likelihood; 2) The teacher-forcing method, which relies on ground truth during training, implies that the model has never encountered its own predictions before testing. RL addresses these discrepancies by bridging the gap between the training and testing phases. Several RL-based approaches for ASR have been proposed [110, 111, 112, 113, 114, 115, 116, 117, 118, 119]. For example, in [110], the authors introduced a RL-based optimization method for the S2S ASR task called Self-critical sequence training (SCST). This method can be conceptualized as a sequential decision model, depicted in Figure 12. The entire encoder-decoder neural network is treated as an agent. At each time step $t$, the current state $s_{t}$ is formed by concatenating the acoustic feature $x_{t}$ and the previous prediction $Y_{t-1}$. The output token serves as the action, updating the generated hypotheses sequence. SCST associates training loss and WER using a WER-related reward function, calculating the reward $r_{t}$ at each token generation step by comparing it with the ground truth sequence $Y^{*}$. SCST uses the test-time beam search algorithm to sample hypotheses for reward normalization, assigning positive weights to high-reward hypotheses that outperform the current test-time system and negative weights to low-reward hypotheses. The framework is illustrated in Figure 12.
In ASR, RL has been primarily proposed to address mismatches between training and testing phases. Two main mismatches leading to potential performance degradation are identified: 1) The commonly used cross-entropy criterion maximizes log-likelihood during training, while performance is evaluated by WER, not log-likelihood; 2) The teacher-forcing method, relying on ground truth during training, implies that the model has never encountered its own predictions before testing. RL bridges the gap between the training and testing phases, addressing these mismatches.
Several RL-based ASR approaches have proposed [110, 111, 112, 113, 114, 115, 116, 117, 118, 119]. For instance, In [110], the authors presented a RL-based optimization method for S2S ASR task called self-critical sequence training (SCST). This can be viewed as a sequential decision model as shown in Fig. 12. The whole encoder-decoder neural network can be viewed as an agent. In each time step $t$, acoustic feature $x_{t}$ and previous $Y_{t-1}$ prediction are concatenated as current state $s_{t}$. The output token is the action at that will update the generated hypotheses sequence. After comparing it with ground truth sequence $Y^{*}$, a reward $r_{t}$ of this time step is calculated. SCST associates the training loss and WER using WER-related reward function, which considers the intermediate reward at each token generation step. Furthermore, SCST utilizes the test-time beam search algorithm to sample a set of hypotheses for reward normalization. As a result, the high-reward hypotheses that outperform the current test-time system are given positive weights, while the low-reward hypotheses are given negative weights.
The entire encoder-decoder neural network is treated as an agent. At each time step $t$, the current state $s_{t}$ is formed by concatenating the acoustic feature $x_{t}$ and the previous prediction $Y_{t-1}$. The output token serves as the action, updating the generated hypotheses sequence. SCST associates training loss and WER using a WER-related reward function, calculating the reward $r_{t}$ at each token generation step by comparing it with the ground truth sequence $Y^{*}$. SCST uses the test-time beam search algorithm to sample hypotheses for reward normalization, assigning positive weights to high-reward hypotheses that outperform the current test-time system and negative weights to low-reward hypotheses. The framework is illustrated in Figure 12.
Figure 12: SCST sequential decision model of ASR [110].
@@
In [111], the authors developed a RL framework for speech recognition systems using the policy gradient method. They introduced a RL method within this framework, incorporating user feedback through hypothesis selection. Tjandra et al. [112, 113] also employed policy gradient RL to train a S2S ASR model. In [114], the authors constructed a generic RL-based AutoML system. This system automatically optimizes per-layer compression ratios for a SOTA attention-based end-to-end ASR model, which consists of multiple LSTM layers. The compression method employed in this work is singular value decomposition (SVD) low-rank matrix factorization. The authors improved this approach by combining iterative compression with AutoML-based rank searching, achieving over 5 x ASR compression without degrading the WER [115]. Shen et al. [116] suggested employing RL to optimize a speech enhancement model based on recognition results, aiming to directly enhance ASR performance. AutoML-based low-rank factorization (LRF) achieves up to 3.7× speedup. In the shade of this, Mehrotra et al. in their work [115] propose an iterative AutoML-based LRF that employs RL for the iterative search, surpassing 5× compression without degrading WERs, advancing ASR. Table 8 summarizes the recent RL-based ASR techniques.
@@
Table 8: Summary of recent proposed works in RL-based ASR. The symbol ($\uparrow$) result increase, whereas ($\downarrow$) signifies result decrease. In cases where multiple scenarios are examined, only the top-performing outcome is mentioned.
Scheme
Model-based
ASR Tasks
RL technique
Metric and result
[110]
S2S Conformer
Improve ASR
Policy gradient

### v2

@@ 3.1.2 Language domain
CTC is an architecture and principle commonly used in S2S tasks (Figure 5 (c)), such as ASR. It enables alignment-free training by introducing a blank symbol and allowing variable-length alignments between input and output sequences. During training, the model learns to align the input sequence with the target sequence, and the blank symbol accounts for multiple possible alignments. CTC is particularly effective in tasks with variable-length outputs, making it well-suited for applications like speech recognition where the duration of spoken words may vary. This latter has been used in many ASR schemes, for example, Deng et al. [54] presents the innovative pretrained Transformer S2S ASR architecture, which integrates self-supervised pretraining techniques for comprehensive end-to-end ASR. Employing a hybrid CTC/attention model, it maximizes the potential of pretrained AM and LM. The inclusion of a CTC branch aids in the encoder’s convergence during training and considers all potential time boundaries in beam searching. The encoder is initiated with wav2vec2.0, and the introduction of a one-cross decoder mitigates reliance on acoustic representations, enabling initialization with pretrained DistilGPT2 and overcoming the constraint of conditioning on acoustic features.
@@ 3.2.1 Acoustic domain
Schneider et al. [82] explored unsupervised pre-training for speech recognition using wav2vec model on large unlabeled audio data. The learned representations enhanced AM training with a simple CNN optimized through noise contrastive binary classification. In [83], a source filter warping data augmentation strategy is proposed to enhance the robustness of children’s speech ASR. The authors constructed an end-to-end acoustic model using the XLS-R wav2vec 2.0 model, pre-trained in a self-supervised manner on extensive cross-lingual corpora of adult speech. The work proposed in [84] introduces a multi-dialect acoustic model employing soft-parameter-sharing multi-task learning, a transductive DTL subcategory, within the Transformer architecture. Auxiliary cross-attentions aid dialect ID recognition, providing dialect information. Adaptive cross-entropy loss automatically balances multi-task learning. Experimental results demonstrate a significant reduction in error rates compared to various single- and multi-task models on multi-dialect speech recognition and dialect ID recognition tasks. Similarly, in the realm of computer vision, CNNs models like ConvNeXt have outperformed cutting-edge Transformers, partly due to the integration of depthwise separable convolutions (DSC). DSC, which approximates regular convolutions, enhances the efficiency of CNNs in terms of time and memory usage without compromising accuracy—in some cases, even enhancing it. The study [85] introduces DSC into the pre-trained audio model family for audio classification on AudioSet (target task), demonstrating its advantages in balancing accuracy and model size. Xin et al. [86] introduce an audio pyramid Transformer with an attention tree structure, with four branches, to reduce computational complexity in fine-grained audio spectrogram processing. It proposes a DA transfer learning approach for weakly supervised sound ED, a sub-field of ASR, enhancing localization performance by aligning feature distributions between frame and clip domains with a DA detection loss.
@@ 3.2.2 Language domain
The methodology is founded on the utilization of the BERT model [87], which involves pretraining language models and demonstrates improved performance across various downstream tasks. DTL approaches for language models, specifically employed in the domain of voice recognition, are referred to as LM adaptation. These approaches aim to bridge the gap between the source distribution $\mathbb{D}_{S}$ and the target distribution $\mathbb{D}_{T}$. Song et al. [88] present L2RS approach, which relies on two main components: (i) utilizing diverse textual data from SOTA NLP models, such as BERT, and (ii) automatically determining their weights to rescore the N-best lists for ASR systems.
Recent advancements in S2S models have shown promising results for training monolingual ASR systems. The CTC and encoder-decoder models are two popular architectures for end-to-end ASR. Additionally, joint training of these architectures in a multi-task hybrid approach has been explored, demonstrating improved overall performance. For instance, the architecture illustrated in Figure 5 (a) comprises S2S layers. The encoder network of S2S consists of a series of RNNs that generate embedding vectors, while the RNN decoder utilizes these vectors to produce final results. The RNN also benefits from prior predictions ($P_{i},i=0,\dots,n$), enhancing the accuracy of subsequent predictions. Moving on, a novel DTL-based approach that enhances end-to-end speech recognition has been proposed in [89]. The novelty lies in applying DTL through multilingual training and multi-task learning at two levels. The initial stage utilizes non-negative matrix factorization, instead of a bottleneck layer, and multilingual training for high-level feature extraction. The subsequent stage employs joint CTC-attention models on these features, where the CTC was transferred to the target attention-based model. The scheme demonstrated superior performance on TIMIT but requires testing on high-resource data. Further optimization is needed for standard end-to-end training. In addition, integrating both AM and LM methodologies has the potential to enhance or construct an effective DTL-based DTL model, as demonstrated in [46, 90, 91, 92, 93, 94].
@@
Scheme Based on Speech recognition task ($\mathbb{T}_{T}$) AM/LM Adaptation Result with metric
[85] ConvNeXt-Tiny Audio classification AM DA mAP= 0.471
[86] APT Sound event detection AM DA F1= 79.6%
[49] BERT (Jasper) Speech-to-text LM DTL WER= 14%
[54] DistilGPT2 Improve ASR Both Fine-tuning CER= 4.6%
[94] XLRS Wave2vec Improve ASR in low resource language Both Fine-tuning 5.6% WER $\downarrow$
[83] XLRS Wave2vec Improve ASR for children’s speech AM DTL WER = 4.86%
@@
[84] Transformer Multi-dialect model aids recognizing diverse speech dialects effectively AM Multi-task learning Acc= 100%
[95] S2S Enhancing the existing multilingual S2S model. LM DTL 4%CER $\downarrow$ 6% WER
[96] PaSST Audio tagging and event detection AM Fine-tuning F1= 64.85%
@@
[97] ARoBERT Spoken language understanding LM Fine-tuning F1-score=92.56%
@@ 3.3 FL-based ASR
  • Horizontal federated learning (HFL): In HFL, clients train a shared global model using their respective datasets, characterized by the same feature space but different sample spaces. Each client utilizes a local AI model, and their updates are aggregated by a central server without exposing raw data. The HFL training process involves: (1) initialization, (2) local training, (3) encryption of gradients, (4) secure aggregation, and (5) global model parameter updates. The objective function minimizes a global loss across all parties’ datasets [98].
[…]
    VFL: vertical federated learning (VFL) trains models on datasets sharing the same sample space but having different feature spaces. Through entity data alignment (EDA) and encrypted model trained (EMT), VFL allows clients to cooperatively train models without sharing raw data. The training process involves the same steps as HFL [98].
@@
FL presents a paradigm shift in AI training, promoting collaboration while respecting data privacy. The working principle of HFL, and VFL is depicted on Figure 9, they cater to various data distribution scenarios, offering flexible solutions for decentralized and secure ML. The application of these FL frameworks extends across diverse domains, promising improved model accuracy and privacy preservation. The first work introducing FL in ASR is presented in [99]. The authors introduced a FL platform that is easily generalizable, incorporating hierarchical optimization and a gradient selection algorithm to enhance training time and SR performance. Guliani et al. [100] proposed a strategy to compensate non-identically distributed (non-IID) data in federated training of ASR systems. The proposed strategy involved random client data sampling, which resulted in a cost-quality trade-off. Zhu et al. [101] addressed also FL-based ASR in non-IID scenarios with personalized FL. They introduced two approaches: adapting personalization layer-based FL for ASR, involving local layers for personalized models, and proposing decoupled federated learning (DecoupleFL). DecoupleFL reduces computation on clients by shifting the computation burden to the server. Additionally, it communicates secure high-level features instead of model parameters, reducing communication costs, particularly for large models. In [102], the authors proposed a client-adaptive federated training scheme to mitigate data heterogeneity when training ASR models. Nguyen et al. [103] used FL to train an ASR model based on a wav2vec 2.0 model pre-trained by self supervision. Yang et al. [104] proposed a decentralized feature extraction approach in federated learning. This approach is built upon a quantum CNN (QCNN) composed of a quantum circuit encoder for feature extraction, and an RNN based end-to-end AM. This framework takes advantage of the quantum learning progress to secure models and to avoid privacy leakage attacks. Gao et al. [105] tackled a challenging and realistic ASR federated experimental setup with clients having heterogeneous data distributions, featuring thousands of different speakers, acoustic environments, and noises. Their empirical study focused on attention-based S2S end-to-end ASR models, evaluating three aggregation weighting strategies: standard FedAvg, loss-based aggregation, and a novel WER-based aggregation. Table 6 summarizes the most recent FL-based ASR techniques.
@@
Scheme Based on Speech recognition task FL technique Metric and result
[99] S2S Improve ASR FedAvg WER = 6%
[100] End-to-end RNN-T ASR on non-IDD data FedAvg WER=
[101] CNN+Transformer extractor ASR on non-IDD data DecoupleFL 2.3- 3.4% WER $\downarrow$ compared with FedAvg
[102] LSTM ASR on non-IDD data CAFT WER = 15.13%
[103] wav2vec 2.0 Improve ASR FedAvg WER= 10.92%
@@
[104] QCNN and RNN Improve privacy-preservation in ASR VFL Accuracy = 95.12%
[105] S2S ASR on heterogeneous data distributions FedAvg WER= 19.98-23.86%
[106] S2S ASR on private and unlabelled user data. FedNST WER= 22.5%
[107] Wav2vec 2.0 and Whisper ASR & KWS for child exploitation settings FedAvg WER = 11-25%
[108] Kaldi and backoff n-gram Improve privacy-preservation in ASR Merging models WER= 17.7%
[109] TDNN Improve privacy-preservation in ASR Aggregation EER = 1-2%.
[110] Non-Streaming & Streaming Conformer Reduce client ASR model size Federated Dropout 6-22% $\downarrow$ Client size model; 34-3% WER $\downarrow$
3.4 DRL-based ASR
@@
DRL is a ML paradigm where an agent learns optimal decision-making by interacting with an environment. The agent receives feedback in the form of rewards or penalties, adapting its behavior to maximize cumulative reward over time through a trial-and-error process. DRL involves several key concepts, as defined in the following terms: Environment model, serving as a representation of contextual dynamics; State ($s$), denoting the current situation perceived by the agent; Observation ($o$), a subset of the state directly perceived by the agent; Action ($a$), the decision made by the agent in response to the environment; Policy ($\pi$), describing how the agent converts environmental conditions into actions; Agent, the entity making decisions based on current states and past experiences; Reward, numerical values assigned by the environment to the agent based on state-action interactions. Figure 10 illustrates the principle of DRL.
Figure 10: DRL principle.
@@
Markov decision process (MDP) is a fundamental framework for dynamic and stochastic decision-making, characterized by state space $S$, action space $\mathbb{A}$, transition probabilities $\mathbb{P}$, and a reward function $R$. The primary objective in an MDP is to identify an optimal policy $\pi^{*}$ maximizing the expected discounted total reward over time, expressed as:
$\pi^{*}=\max_{\pi}\mathbb{E}_{\pi}[\sum_{t=0}^{T}\gamma^{t}r_{t}(s_{t},\pi(s_{t}))],$ (4)
where $\gamma$ is the discount factor. MDPs find extensive applications in addressing uncertainties in intelligent systems within dynamic wireless environments, including spectrum management, cognitive radios, and wireless security.
@@
In the field of ASR, DRL has primarily been proposed to tackle discrepancies between the training and testing phases. Two main discrepancies leading to potential performance deterioration have been identified: 1) The conventional use of the cross-entropy criterion maximizes log-likelihood during training, while performance is assessed by WER, not log-likelihood; 2) The teacher-forcing method, which relies on ground truth during training, implies that the model has never encountered its own predictions before testing. DRL addresses these discrepancies by bridging the gap between the training and testing phases. Several DRL-based approaches for ASR have been proposed [111, 112, 113, 114, 115, 116, 117, 118, 119, 120]. For example, in [111], the authors introduced a DRL-based optimization method for the S2S ASR task called self-critical sequence training (SCST). This method can be conceptualized as a sequential decision model, depicted in Figure 11. The entire encoder-decoder neural network is treated as an agent. At each time step $t$, the current state $s_{t}$ is formed by concatenating the acoustic feature $x_{t}$ and the previous prediction $Y_{t-1}$. The output token serves as the action, updating the generated hypotheses sequence. SCST associates training loss and WER using a WER-related reward function, calculating the reward $r_{t}$ at each token generation step by comparing it with the ground truth sequence $Y^{*}$. SCST uses the test-time beam search algorithm to sample hypotheses for reward normalization, assigning positive weights to high-reward hypotheses that outperform the current test-time system and negative weights to low-reward hypotheses. The framework is illustrated in Figure 11.
Figure 11: Example of sequential decision model of DRL-based ASR [111].
@@
In [112], the authors developed a DRL framework for speech recognition systems using the policy gradient method. They introduced a DRL method within this framework, incorporating user feedback through hypothesis selection. Tjandra et al. [113, 114] also employed policy gradient DRL to train a S2S ASR model. In [115], the authors constructed a generic DRL-based AutoML system. This system automatically optimizes per-layer compression ratios for a SOTA attention-based end-to-end ASR model, which consists of multiple LSTM layers. The compression method employed in this work is singular value decomposition (SVD) low-rank matrix factorization. The authors improved this approach by combining iterative compression with AutoML-based rank searching, achieving over 5 x ASR compression without degrading the WER [116]. Shen et al. [117] suggested employing DRL to optimize a SE model based on recognition results, aiming to directly enhance ASR performance. AutoML-based low-rank factorization (LRF) achieves up to 3.7× speedup. In the shade of this, Mehrotra et al. in their work [116] propose an iterative AutoML-based LRF that employs DRL for the iterative search, surpassing 5× compression without degrading WERs, advancing ASR. Table 7 summarizes the recent DRL-based ASR techniques.
@@
Table 7: Summary of recent proposed works in DRL-based ASR. The symbol ($\uparrow$) result increase, whereas ($\downarrow$) signifies result decrease. In cases where multiple scenarios are examined, only the top-performing outcome is mentioned.
Scheme Model-based ASR Tasks DRL technique Metric and result